# An abstract framework for heterogeneous coupling: stability, approximation and applications

### v1

@@ Abstract.
Introducing a coupling framework reminiscent of FETI methods, but here on abstract form, we establish conditions for stability and minimal requirements for well-posedness on the continuous level, as well as conditions on local solvers for the approximation of subproblems. We then discuss stability of the resulting Lagrange multiplier methods and show stability under a mesh conditions between the local discretizations and the mortar space. If this condition is not satisfied we show how a stabilization, acting only on the multiplier can be used to achieve stability. The design of preconditioners of the Schur complement system is discussed in the unstabilized case. Finally we discuss some applications that enter the framework.
@@ 1. Introduction
The common approach in this framework is to discretize the global continuous problem into a possibly huge algebraic system, and then split the latter as a system of coupled algebraic subsystems. In this case the coupling condition often consists in the identification of unknowns in the subsystems corresponding to the same global unknown. An alternative approach, which is lately gaining increasing interest, is, instead, to decompose the original problem already at the continuous/infinite dimensional level, thus obtaining a system of coupled infinite dimensional subproblems, to be successively discretized. This is the approach that underlies a number of non conforming domain decomposition methods, such as the mortar method [4, 1, 31], the three fields domain decomposition method [11, 7], or the domain decomposition method based on local Dirichlet–Neumann maps proposed in [25], which all start from a domain decomposition of the continuous problem.
@@
A key feature of this approach is that the subproblems can use different models and be discretized independently of each other using bespoke solvers that communicate only through the coupling conditions. This offers the possibility of resorting to independent implementations for the solvers of the different subproblems, an appealing feature, particularly in the framework of multiphysics and multiscale simulations, where highly specialized and optimized subproblem solvers might be available, and where the whole coupled problem might be excessively complex. Indeed, in this framework, we see a growing interest around the different issues related to “code coupling” [22, 21, 24], and the development of different computational environments aimed at enabling the interaction of existing solvers, individually simulating different portions of a larger complex physical problems (see, for instance, [28, 12, 30]).
@@
The aim of this paper is to provide a theoretical foundation for such approaches where heterogeneous large scale problems are solved using different highly optimized codes to handle the different subproblems. Therefore we study, in an abstract framework, with problem, subproblems and coupling that are not a priori associated to a physical space and model, a strategy which is agnostic to the different solvers, assuming only that they produce a sufficiently accurate approximation of their target quantities. Following an approach inspired by the widely used Finite Element Tearing and Interconnecting (FETI) method [20, 18, 19, 23], by formally solving the local problems, up to a possible non-trivial, but finite dimensional, kernel of the local operator, the system is reduced to a problem where only the unknowns playing the role of Lagrangian multipliers, enforcing the coupling between subproblems, are to be computed. While the FETI method was developed in the framework of conforming domain decomposition for PDEs, we believe a large class of other problems, including multiscale models [17], FEM/BEM coupling [6], or PDEs set on lower dimensional manifolds connected in a network, can benefit from it and from the related techniques developed to handle the resulting algebraic system.
@@
Our abstract starting point allows us to remain agnostic as to the local approximation strategy to the furthest possible extent: discretization of the local subproblem consists, in our abstract famework, in selecting a non-specified “local solver”. This important feature of our approach allows the coupling of highly optimized codes with solvers that use different methods, for instance finite difference methods, finite element methods, spectral methods or boundary element methods. By carrying out our analysis in an abstract framework, independent of the discretization methodology, we aim to establish the weakest conditions that such solvers must satisfy in order to guarantee the well posedness of the resulting global method. To establish such minimal conditions we analyze the properties of the reduced coupling system, for which well-posedness is proven at the continuous level. This continuous framework is then exploited when local discretizations are introduced, which we treat as “black boxes”, and conditions are proposed that are sufficient (but not necessary) to establish uniform stability of the discretization. To cater for the situation where such conditions, which implicitely enforce a compatibility condition between local solvers and multiplier discretization, can not be satisfied or lead to inconvenient discretizations, we introduce stabilization methods in the spirit of [10], generalizing the ideas of [15, 14]. We establish optimal error estimates for the discretization and we discuss the design of preconditioners for the unstabilized reduced system. Finally we discuss some different examples that enter the framework.
@@ 2. Problem setting
[…]
@@
[…]
@@
It will be convenient, in the following, to write Problem 2.2 in weak form as: find $\lambda\in\Lambda$ and $z\in Z$ such that
$s(\lambda,\mu)+b(z,\mu)$ $=$ $\langle g,\mu\rangle,\qquad\forall\mu_{\delta}\in\Lambda_{\delta},$
@@
This new saddle point problem, defined already at the continuous level, is the starting point of our abstract coupling algorithm. We start by verifying the well posedness of such a problem. We have the following Theorem.
@@
$s(\lambda,\mu)=\langle A^{+}B^{T}\lambda,B^{T}\mu\rangle\lesssim\|A^{+}B^{T}\lambda\|_{V}\|A^{+}B^{T}\mu\|_{V}\lesssim\|\lambda\|_{\Lambda}\|\mu\|_{\Lambda}.$
@@
Problem 2.2 admits a unique solution $(\lambda,z)$ such that, letting $\mathsf{u}$ denote the solution of (2.1), we have
@@ 3. Discretization
We discretize Problem 2.2 by a non conforming Galerkin discretization, involving both approximating the multiplier space $\Lambda$ with a finite dimensional space $\Lambda_{\delta}$, and numerically evaluating the pseudo inverses $A_{k}^{+}$. More precisely, for all $k$ we consider approximate linear solvers $A^{+}_{h,k}:V^{\prime}\to C_{k}$, that we will treat as a black-boxes. We then let
@@
As we want to treat them as black boxes, we want the assumptions on the approximate pseudo inverses to be as weak as possible. We want to understand under which minimal conditions Problem 3.1 is an approximation to Problem 2.2. To this aim it is only natural to require that the black box approximate linear solvers $A^{+}_{h,k}$ are approximations of the continuous pseudo inverses $A^{+}_{k}$, at least when the data have some extra smoothness.
To quantify, in our abstract setting, the concept of “extra smoothness”, we introduce infinite dimensional subspaces of $V_{k}$ and $\Lambda$, playing the role of higher regularity spaces. More precisely we let $W^{r}_{k}\subseteq V_{k}$ and $\Xi^{r}\subseteq\Lambda$, $r\geq 0$, denote a scale of subspaces of $V_{k}$ and $\Lambda$ respectively, such that $W^{0}_{k}=V_{k}$ and $\Xi^{0}=\Lambda$, and such that, for $t<r$, $\Xi^{r}\subset\Xi^{t}$ and $W^{r}_{k}\subset W_{k}^{t}$. We assume that, in this smoothness scale, $A^{+}B^{T}$ is smoothness preserving up to a certain range $r_{0}$, that is, that $\lambda\in\Xi^{r}$ implies that $A^{+}B^{T}\lambda\in W^{r}=\prod_{k}W^{r}_{k}$, and
(3.1) $\|A^{+}B^{T}\lambda\|_{W^{r}}\leq C_{r}\|\lambda\|_{\Xi^{r}},\qquad 0\leq r\leq r_{0}.$
@@
The analysis below also requires the following inverse inequality to hold true in $\Lambda_{\delta}$.
@@
We start by observing that, for $r>0$ sufficiently small, for $\lambda\in\Lambda$ and $\mu_{\delta}\in\Lambda_{\delta}$, we can write
@@
which proves the continuity. Now, note that if $\lambda=\lambda_{\delta}\in\Lambda_{\delta}$ in (3.3), then by Assumption 3.5 we have, for some $r>0$,
@@
Under Assumptions 3.3, 3.5 and 3.7, $u-z\in W^{r}$, $r<R$ implies
@@
$\|\lambda_{\delta}-\xi_{\delta}\|_{\Lambda}\lesssim s_{h}(\lambda_{\delta}-\xi_{\delta},\lambda_{\delta}-\xi_{\delta})\lesssim s_{h}(\lambda-\xi_{\delta},\lambda_{\delta}-\xi_{\delta})$
@@
where we used Assumption 3.2 and (2.6) for $I$ followed by the boundedness of $A^{+}_{h}$ and the bound $\|B^{T}(\lambda-\lambda_{\delta})\|_{V^{\prime}}\lesssim\|\lambda-\lambda_{\delta}\|_{\Lambda}$ for $II$. ∎
@@
Under Assumptions 3.2, 3.4, 3.5 and 3.7, if $u\in W^{r}$, $r\leq R$ and $\lambda\in\Xi^{t}$, $t\leq m$, setting $u_{h}=A^{+}_{h}(f+B^{T}\lambda_{\delta})+z^{*}$ we have that
@@
Note that it is not the smoothness of $u$ that comes into play in Theorem 3.8, but the smoothness of $u-z$.
@@ 3.1. Including stabilization
Consider now a situation where the condition $h/\delta<C_{\Lambda}$ with $C_{\Lambda}$ sufficiently small, given by Lemma 3.5, does not hold, but where we can not (or do not want to) take a smaller $h$ or a larger $\delta$. Assume however that we know that (3.2) holds, with a coarser mesh-size $\tilde{\delta}$ for which $h/\tilde{\delta}<C_{\Lambda}$, for some auxiliary space $\widetilde{\Lambda}_{\delta}$, not necessarily related to $\Lambda_{\delta}$, but with similar asymptotic properties of approximation. This can be exploited to design a stabilization term and retrieve the well posedness of the discrete problem.
To this aim we let $\tilde{\pi}_{\delta}:\Lambda\to\tilde{\Lambda}_{\delta}$ denote a bounded projection on $\tilde{\Lambda}_{\delta}$. Moreover we let $[\cdot,\cdot]_{\delta}:(\Lambda_{\delta}+\tilde{\Lambda}_{\delta})\times(\Lambda_{\delta}+\tilde{\Lambda}_{\delta})\to\mathbb{R}$ denote a continuous bilinear form satisfying
$[\lambda_{\delta}-\tilde{\pi}_{\delta}\lambda_{\delta},\lambda_{\delta}-\tilde{\pi}_{\delta}\lambda_{\delta}]_{\delta}\gtrsim\|\lambda_{\delta}-\tilde{\pi}_{\delta}\lambda_{\delta}\|^{2}_{\Lambda},\qquad\text{ for all }\lambda_{\delta}\in\Lambda_{\delta},$
@@
Assume that the black box approximate solver $A_{h}^{+}$ satisfies Assumptions 3.3 and 3.7, and that the space $\tilde{\Lambda}_{\delta}$ satisfies Assumption 3.5. Then, there exists $\gamma_{0}$ such that, provided $\gamma>\gamma_{0}$, Problem 3.12 is well posed, and the following error bound holds for its solution:
@@
[…]
In the first term we first add and subtract $-BA^{+}B^{T}\tilde{\pi}_{\delta}\lambda_{\delta}$ on the left slot to get
[…]
and then, letting $c_{s}$ be the coercivity constant of $s(\cdot,\cdot)$, we bound from below the second term on the right hand side by adding and subtracting $-BA^{+}B^{T}\lambda_{\delta}$, which allows us to write
$BA^{+}B^{T}\tilde{\pi}_{\delta}\lambda_{\delta},\lambda_{\delta}\right>=\left<BA^{+}B^{T}(\tilde{\pi}_{\delta}-I)\lambda_{\delta},\lambda_{\delta}\right>+s(\lambda_{\delta},\lambda_{\delta})\\ \geq\left<BA^{+}B^{T}(\tilde{\pi}_{\delta}-I)\lambda_{\delta},\lambda_{\delta}\right>+c_{s}\|\lambda_{\delta}\|_{\Lambda}^{2}.$
@@
$BA^{+}B^{T}(\tilde{\pi}_{\delta}-I)\lambda_{\delta},\lambda_{\delta}\right>\leq C\|\tilde{\pi}_{\delta}\lambda_{\delta}-\lambda_{\delta}\|_{\Lambda}\|\lambda_{\delta}\|_{\Lambda}\leq C\sqrt{j(\lambda_{\delta},\lambda_{\delta})}\|\lambda_{\delta}\|_{\Lambda}\\ \leq c_{1}\epsilon^{-1}j(\lambda_{\delta},\lambda_{\delta})+\frac{1}{4}\epsilon\|\lambda_{\delta}\|_{\Lambda}^{2}.$
@@
$BA_{h}^{+}B^{T}(\tilde{\pi}_{\delta}-I)\lambda_{\delta},\lambda_{\delta}\right>\leq c_{2}\epsilon^{-1}j(\lambda_{\delta},\lambda_{\delta})+\frac{1}{4}\epsilon\|\lambda_{\delta}\|_{\Lambda}^{2}.$
@@
$B(A_{h}^{+}-A)B^{T}\tilde{\pi}_{\delta}\lambda_{\delta},\lambda_{\delta}\right>\leq Ch^{r}\|\tilde{\pi}_{\delta}\lambda_{\delta}\|_{\Xi^{r}}\|\lambda_{\delta}\|_{\Lambda}\leq\frac{h^{r}}{\tilde{\delta}^{r}}\|\tilde{\pi}_{\delta}\lambda_{\delta}\|_{V}\|\lambda_{\delta}\|_{\Lambda}\leq c_{3}\frac{h^{r}}{\tilde{\delta}^{r}}\|\lambda_{\delta}\|_{\Lambda}^{2}.$
@@
Considering this bound in the stability estimate of Problem 3.12 we see that
@@
For the error analysis we proceed as in Theorem 3.8, using now the bound (3.7). Let once again $\xi_{\delta}\in\Lambda_{\delta}$ satisfy (3.4), and let $e_{h}=\lambda_{\delta}-\xi_{\delta}$, with now $\lambda_{\delta}$ solution to Problem 3.12. We have
@@
We note that since $\Lambda_{\delta}$ is finite dimensional, it is often possible to choose $j(\xi_{\delta},\xi_{\delta})$ of a form convenient for computation. Indeed in some cases the projection operator can be eliminated and replaced by some other operator acting directly on $\xi_{\delta}$. This is the case for instance when, in the domain decomposition framework (see Section 5.1), $\widetilde{\Lambda}_{\delta}$ and $\Lambda_{\delta}$ are different finite element spaces of different mesh size, but with the same polynomial order. Then typically $[\xi_{\delta}-\tilde{\pi}_{\delta}\xi_{\delta},\xi_{\delta}-\tilde{\pi}_{\delta}\xi_{\delta}]_{\delta}$ can be replaced by a penalty term acting on the jumps of derivatives of $\xi_{\delta}$ alone, see [14]. In such a framework, computable bilinear forms $[\cdot,\cdot]_{\delta}$ can also be designed by resorting to suitable localization results, see [8]. For an alternative way of stabilizing in an abstract framework, see also [5].

### v2

# An abstract framework for heterogeneous coupling: stability, approximation and preconditioning

@@ Abstract.
We consider heterogeneous coupling problems on an abstract level, establishing fundamental principles of domain decomposition agnostic to the solvers of the local subproblems. Introducing a coupling framework reminiscent of FETI methods, but here on abstract form, we establish conditions for stability and minimal requirements for well-posedness on the continuous level, as well as conditions on local solvers for the approximation of subproblems. We then discuss stability of the resulting Lagrange multiplier methods and show stability under a mesh condition between the local discretizations and the mortar space. If this condition is not satisfied we show how a stabilization, acting only on the multiplier can be used to achieve stability. The design of preconditioners of the Schur complement system is discussed in the unstabilized case. Finally we discuss some applications that enter the framework.
@@ 1. Introduction
The common approach in this framework is to discretize the global continuous problem into a possibly huge algebraic system, and then split the latter as a system of coupled algebraic subsystems. In this case the coupling condition often consists in the identification of unknowns in the subsystems corresponding to the same global unknown. An alternative approach, which is lately gaining increasing interest, is, instead, to decompose the original problem already at the continuous/infinite dimensional level, thus obtaining a system of coupled infinite dimensional subproblems, to be successively discretized. This is the approach that underlies a number of non conforming domain decomposition methods, such as the mortar method [6, 4, 45], the three fields domain decomposition method [20, 12], or the domain decomposition method based on local Dirichlet–Neumann maps proposed in [39].
@@
A key feature of this approach, that we will exploit herein, is that the subproblems can use different models and be discretized independently of each other using bespoke solvers that communicate only through the coupling conditions. This offers the possibility of resorting to independent implementations for the solvers of the different subproblems, an appealing feature, particularly in the framework of multiphysics and multiscale simulations, where highly specialized and optimized subproblem solvers might be available, and where the whole coupled problem might be excessively complex. Indeed, in this context, we see a growing interest around the different issues related to “code coupling” [34, 31, 36], and the development of different computational environments aimed at enabling the interaction of existing solvers, individually simulating different portions of a larger complex physical problems (see, for instance, [41, 21, 43]).
@@
The aim of this paper is to provide a theoretical foundation for such approaches where heterogeneous large scale problems are solved using different highly optimized codes to handle the different subproblems. Therefore we study a strategy, agnostic to the different solvers and subproblems, in an abstract framework. The problem, the subproblems and coupling are not a priori associated to a physical space and model, and therefore a wide range of applications enter the framework. For the local solvers we only assume that they produce a sufficiently accurate approximation of the coupling quantities. Following an approach inspired by the widely used Finite Element Tearing and Interconnecting (FETI) method [29, 27, 28, 35], by formally solving the local problems, up to a possible non-trivial, but finite dimensional, kernel of the local operator, the system is reduced to a problem where only the unknowns enforcing the coupling between subproblems, that is the Lagrangian multipliers, are to be computed. While the FETI method was developed in the framework of conforming domain decomposition for PDEs, the analysis herein applies to a large class of other problems, including multiscale models [25], FEM/BEM coupling [10], or PDEs set on lower dimensional manifolds connected in a network. It also provides a roadmap to techniques for solution of the resulting algebraic system.
@@
Our abstract starting point allows us to remain agnostic as to the local approximation strategy to the furthest possible extent: discretization of the local subproblem consists, in our abstract famework, in selecting a non-specified “local solver”. This important feature of our approach allows the coupling of highly optimized codes with solvers that use different methods, for instance finite difference methods, finite element methods, spectral methods or boundary element methods. By carrying out our analysis in an abstract framework, independent of the discretization methodology, we establish the weakest conditions that such solvers must satisfy in order to guarantee the well posedness of the resulting global method. To establish such minimal conditions we analyze the properties of the reduced coupling system, for which well-posedness is proven at the continuous level. This continuous framework is then exploited when local discretizations are introduced, which we treat as “black boxes”, and conditions are proposed that are sufficient (but not necessary) to establish uniform stability of the discretization. To cater for the situation where such conditions, which implicitely enforce a compatibility condition between local solvers and multiplier discretization, can not be satisfied or lead to inconvenient discretizations, we introduce stabilization methods in the spirit of [18], generalizing the ideas of [23, 22]. We establish optimal error estimates for the discretization and we discuss the design of preconditioners for the unstabilized reduced system. A particularly salient feature of the black box framework is that local solvers based on non-coercive formulations can be preconditioned using alternative coercive formulations. This observation points to a new class of preconditioners for saddle point problems based on local stabilised formulations. The two canonical examples of domain decomposition, Neumann-Neumann and Dirichlet-Dirichlet coupling are used as applications to illustrate the theory.
@@ 2. Problem setting
(2.1) $\mathscr{A}\mathsf{u}=\mathsf{f},\quad\text{ in }\mathscr{V}^{\prime},$
@@
Above and in the following we use the notation $\langle\cdot,\cdot\rangle$ to denote different duality pairings (the spaces paired follow from the context in each case).
@@
Remark 2.2.
Observe that one can also consider a slightly more general form of Problem 2.1, where $g\in\Lambda^{\prime}$ and $C:\Lambda\to\Lambda^{\prime}$ is a positive semi definite linear operator.
Problem 2.3.
Find $u\in V$ and $\lambda\in\Lambda$ such that
(2.5) $Au-B^{T}\lambda$ $=$ $f,$
(2.6) $C\lambda+Bu$ $=$ $g.$
In the PDE context this allows to include Robin type interface coupling conditions between subdomains. The below discussion of well posedness and discretization extends directly to this case.
@@
It will be convenient, in the following, to write Problem 2.4 in weak form as: find $\lambda\in\Lambda$ and $z\in Z$ such that
$s(\lambda,\mu)+b(z,\mu)$ $=$ $\langle g,\mu\rangle,\qquad\forall\mu\in\Lambda,$
@@
This new saddle point problem is the starting point of our abstract coupling algorithm. We start by verifying the well posedness of such a problem. We have the following Theorem.
@@
$s(\lambda,\mu)=\langle A^{+}B^{T}\lambda,B^{T}\mu\rangle\lesssim\|{A^{+}}B^{T}\lambda\|_{V}\|B^{T}\mu\|_{{V^{\prime}}}\lesssim\|\lambda\|_{\Lambda}\|\mu\|_{\Lambda}.$
@@
Problem 2.4 admits a unique solution $(\lambda,z)$ such that, letting $\mathsf{u}$ denote the solution of (2.1), we have
@@ 3. Discretization
We discretize Problem 2.4 by a Galerkin discretization, involving both approximating the multiplier space $\Lambda$ with a finite dimensional space $\Lambda_{\delta}$, and numerically evaluating the pseudo inverses $A_{k}^{+}$. More precisely, for all $k$ we consider approximate linear solvers $A^{+}_{h,k}:V^{\prime}\to C_{k}$, that we will treat as a black-boxes. We then let
@@
As we want to treat the local solvers as black boxes, we want the assumptions on the approximate pseudo inverses to be as weak as possible. We want to understand under which minimal conditions Problem 3.1 is an approximation to Problem 2.4. To this aim it is only natural to require that the black box approximate linear solvers $A^{+}_{h,k}$ are approximations of the continuous pseudo inverses $A^{+}_{k}$, at least when the data have some extra smoothness.
To quantify, in our abstract setting, the concept of “extra smoothness”, we introduce infinite dimensional subspaces of $V_{k}$ and $\Lambda$, playing the role of higher regularity spaces. More precisely we let $W^{r}_{k}\subseteq V_{k}$ and $\Xi^{r}\subseteq\Lambda$, $r\geq 0$, denote a scale of subspaces of $V_{k}$ and $\Lambda$ respectively, such that $W^{0}_{k}=V_{k}$ and $\Xi^{0}=\Lambda$, and such that, for $t<r$, $\Xi^{r}\subset\Xi^{t}$ and $W^{r}_{k}\subset W_{k}^{t}$. We assume that, in this smoothness scale, $A^{+}B^{T}$ is smoothness preserving up to a certain range $R$, that is, that $\lambda\in\Xi^{r}$ implies that $A^{+}B^{T}\lambda\in W^{r}=\prod_{k}W^{r}_{k}$, and
(3.1) $\|A^{+}B^{T}\lambda\|_{W^{r}}\leq C_{r}\|\lambda\|_{\Xi^{r}},\qquad 0\leq r\leq R.$
@@
The analysis below also requires the following inverse inequality to hold true in $\Lambda_{\delta}$, where for the sake of simplicity and without loss of generality we assume that $t\leq R$ ($R$ being the parameter in (3.1)).
@@
We start by observing that for $r$ with $0<r<\min\{t,R\}$, by (3.1) and Assumptions 3.3, for $\lambda\in\Lambda$ and $\mu_{\delta}\in\Lambda_{\delta}$, we can write
@@
which proves the continuity.
Now, note that if $\lambda=\lambda_{\delta}\in\Lambda_{\delta}$ in (3.3), then by Assumption 3.5 we have, for some $r>0$,
@@
Remark 3.7.
The condition appearing in Lemma 3.6 for the coercivity of the Schur complement is similar to that introduced in [2] for the discrete stability of the full mixed Lagrange multiplier problem. We also refer to [17] for a discussion on the well-posedness and approximation of Lagrange multiplier formulations.
@@
Under Assumptions 3.3, 3.5 and 3.8, $u-z\in W^{r}$, $r\leq R$ implies
@@
$\|\lambda_{\delta}-\xi_{\delta}\|_{\Lambda}^{2}\lesssim s_{h}(\lambda_{\delta}-\xi_{\delta},\lambda_{\delta}-\xi_{\delta})\lesssim s_{h}(\lambda-\xi_{\delta},\lambda_{\delta}-\xi_{\delta})$
@@
where we used Assumption 3.2 and (2.8) for II followed by the boundedness of $A^{+}_{h}$ and the bound $\|B^{T}(\lambda-\lambda_{\delta})\|_{V^{\prime}}\lesssim\|\lambda-\lambda_{\delta}\|_{\Lambda}$ for $II$. ∎
@@
Under Assumptions 3.2, 3.4, 3.5 and 3.8, if $u\in W^{r}$, $r\leq R$ and $\lambda\in\Xi^{t}$, $t\leq m$, setting $u_{h}=A^{+}_{h}(f+B^{T}\lambda_{\delta})+z^{*}$ we have that
@@
Note that it is not the smoothness of $u$ that comes into play in Theorem 3.9, but the smoothness of uzu-z.
Remark 3.13.
Our framework requires the bilinear form $a$ to be coercive on $(\ker A_{k})^{\perp}$. On the other hand, it often happens that $a$ is only semi definite on $(\ker A_{k})^{\perp}$ (for instance when the local problems are themselves saddle point problems), satisfying inf-sup conditions of the form
$\inf_{v\in(\ker A_{k})^{\perp}}\sup_{w\in(\ker A_{k})^{\perp}}\frac{a_{k}(v,w)}{\|v\|_{V_{k}}\|w\|_{V_{k}}}\gtrsim 1,\qquad\inf_{w\in(\ker A_{k})^{\perp}}\sup_{v\in(\ker A_{k})^{\perp}}\frac{a_{k}(v,w)}{\|v\|_{V_{k}}\|w\|_{V_{k}}}\gtrsim 1.$
However, following the approach of [3], it is always possible to write down an equivalent coercive form of the local problem falling in our framework, by adding a residual term measured in the $V_{k}^{\prime}$ norm. This approach allows us to extend our analysis to more general situations. By treating local numerical solvers as black boxes, we avoid dependence on the specific formulation of the local problem. Consequently, these solvers can be designed according to the original (non-coercive) formulation.
@@ 3.1. Including stabilization
Consider a situation where the condition $h/\delta<C_{\Lambda}$ as given by Lemma 3.6, does not hold, but where we can not (or do not want to) reduce $h$ or increase $\delta$. Assume however that we know (3.2) holds for a coarser mesh-size $\tilde{\delta}$ where $h/\tilde{\delta}<C_{\Lambda}$, for some auxiliary space $\widetilde{\Lambda}_{\delta}$ that is not necessarily related to $\Lambda_{\delta}$ but has similar asymptotic approximation properties. This can be leveraged to design a stabilization term and ensure the well-posedness of the discrete problem.
To this aim we let $\tilde{\pi}_{\delta}:\Lambda\to\tilde{\Lambda}_{\delta}$ denote a bounded projection on $\tilde{\Lambda}_{\delta}$. Moreover we let $\cdot,\cdot]_{\delta}:(\Lambda_{\delta}+\tilde{\Lambda}_{\delta})\times(\Lambda_{\delta}+\tilde{\Lambda}_{\delta})\to\mathbb{R}$ denote a continuous bilinear form satisfying
$\lambda_{\delta}-\tilde{\pi}_{\delta}\lambda_{\delta},\lambda_{\delta}-\tilde{\pi}_{\delta}\lambda_{\delta}]_{\delta}\gtrsim\|\lambda_{\delta}-\tilde{\pi}_{\delta}\lambda_{\delta}\|^{2}_{\Lambda},\qquad\text{ for all }\lambda_{\delta}\in\Lambda_{\delta},$
@@
Assume that the black box approximate solver $A_{h}^{+}$ satisfies Assumptions 3.3 and 3.8, and that the space $\tilde{\Lambda}_{\delta}$ satisfies Assumption 3.5. Then, there exists $\gamma_{0}$ such that, provided $\gamma>\gamma_{0}$, Problem 3.14 is well posed, and the following error bound holds for its solution:
@@
$s_{h}(\lambda_{\delta},\lambda_{\delta})=\left<BA_{h}^{+}B^{T}\lambda_{\delta},\lambda_{\delta}\right>=\left<BA_{h}^{+}B^{T}\tilde{\pi}_{\delta}\lambda_{\delta},\lambda_{\delta}\right>+\left<BA_{h}^{+}B^{T}(I-\tilde{\pi}_{\delta})\lambda_{\delta},\lambda_{\delta}\right$>.
In the first term we first add and subtract $BA^{+}B^{T}\tilde{\pi}_{\delta}\lambda_{\delta}$ on the left slot to get
$\left<BA_{h}^{+}B^{T}\tilde{\pi}_{\delta}\lambda_{\delta},\lambda_{\delta}\right>=\left<B(A_{h}^{+}-A)B^{T}\tilde{\pi}_{\delta}\lambda_{\delta},\lambda_{\delta}\right>+\left<BA^{+}B^{T}\tilde{\pi}_{\delta}\lambda_{\delta},\lambda_{\delta}\right$>,
and then, letting $c_{s}$ be the coercivity constant of $s(\cdot,\cdot)$, we bound from below the second term on the right hand side by adding and subtracting $BA^{+}B^{T}\lambda_{\delta}$, which allows us to write
$\left<BA^{+}B^{T}\tilde{\pi}_{\delta}\lambda_{\delta},\lambda_{\delta}\right>=\left<BA^{+}B^{T}(\tilde{\pi}_{\delta}-I)\lambda_{\delta},\lambda_{\delta}\right>+s(\lambda_{\delta},\lambda_{\delta})\\ \geq\left<BA^{+}B^{T}(\tilde{\pi}_{\delta}-I)\lambda_{\delta},\lambda_{\delta}\right>+c_{s}\|\lambda_{\delta}\|_{\Lambda}^{2}.$
@@
$\left<BA^{+}B^{T}(\tilde{\pi}_{\delta}-I)\lambda_{\delta},\lambda_{\delta}\right>\leq C\|\tilde{\pi}_{\delta}\lambda_{\delta}-\lambda_{\delta}\|_{\Lambda}\|\lambda_{\delta}\|_{\Lambda}\leq C\sqrt{j(\lambda_{\delta},\lambda_{\delta})}\|\lambda_{\delta}\|_{\Lambda}\\ \leq c_{1}\epsilon^{-1}j(\lambda_{\delta},\lambda_{\delta})+\frac{1}{4}\epsilon\|\lambda_{\delta}\|_{\Lambda}^{2}.$
@@
$\left<BA_{h}^{+}B^{T}(\tilde{\pi}_{\delta}-I)\lambda_{\delta},\lambda_{\delta}\right>\leq c_{2}\epsilon^{-1}j(\lambda_{\delta},\lambda_{\delta})+\frac{1}{4}\epsilon\|\lambda_{\delta}\|_{\Lambda}^{2}.$
@@
$\left<B(A_{h}^{+}-A)B^{T}\tilde{\pi}_{\delta}\lambda_{\delta},\lambda_{\delta}\right>\leq Ch^{r}\|\tilde{\pi}_{\delta}\lambda_{\delta}\|_{\Xi^{r}}\|\lambda_{\delta}\|_{\Lambda}\leq\frac{h^{r}}{\tilde{\delta}^{r}}\|\tilde{\pi}_{\delta}\lambda_{\delta}\|_{V}\|\lambda_{\delta}\|_{\Lambda}\leq c_{3}\frac{h^{r}}{\tilde{\delta}^{r}}\|\lambda_{\delta}\|_{\Lambda}^{2}.$
@@
Considering this bound in the stability estimate of Problem 3.14 we see that
@@
For the error analysis we proceed as in Theorem 3.9, using now the bound (3.7). Let once again $\xi_{\delta}\in\Lambda_{\delta}$ satisfy (3.4), and let $e_{h}=\lambda_{\delta}-\xi_{\delta}$, with now $\lambda_{\delta}$ solution to Problem 3.14. We have
@@
We note that since $\Lambda_{\delta}$ is finite dimensional, it is often possible to choose $j(\xi_{\delta},\xi_{\delta})$ on a form convenient for computation. Indeed in some cases the projection operator can be eliminated and replaced by some other operator acting directly on $\xi_{\delta}$. This is the case for instance when, in the domain decomposition framework (see Section 5.1), $\widetilde{\Lambda}_{\delta}$ and $\Lambda_{\delta}$ are different finite element spaces of different mesh size, but with the same polynomial order. Then typically $\xi_{\delta}-\tilde{\pi}_{\delta}\xi_{\delta},\xi_{\delta}-\tilde{\pi}_{\delta}\xi_{\delta}]_{\delta}$ can be replaced by a penalty term acting on the jumps of derivatives of $\xi_{\delta}$ alone, see [22]. In such a framework, computable bilinear forms $\cdot,\cdot]_{\delta}$ can also be designed by resorting to suitable localization results, see [13]. For an alternative way of stabilizing in an abstract framework, see also [9].